LM= 3x, MN= 6, and LN= 5x, what is LM


1
Expert's answer
2020-10-26T20:25:27-0400

If LM = 3x, MN = 6, and LN = 5x, then as LM + MN = LN, so:

3x + 6 = 5x,

2x = 6,

x = 3,

LM = 3×3 = 9.